OpenStudy (anonymous):

Find the Domain of the following question. I know that the denominator has to be more than 0, and the radical cant be negative, but how do I figure this out? See reply for the function.

OpenStudy (anonymous):

\(\ \Huge \frac{1}{\sqrt[4]{x^2-5x}}. \)

OpenStudy (anonymous):

You have to solve the inequation: x^2 -5 > 0.

OpenStudy (anonymous):

What about \(\ \Huge \sqrt[4]{} ?\)

OpenStudy (anonymous):

The solution will be positive included the root

OpenStudy (anonymous):

How do you solve \(\ \Huge x^2-5>0?\)

OpenStudy (anonymous):

x^2 -5x > 0 ; x( x-5) > 0 ; x> 5 or x>0